OpenStudy (anonymous):

of the teachers in a school 3/4 are women. of the women teacher in the school 5/6 are married. work out the fraction of teachers who are married. @Abhisar

OpenStudy (anonymous):

11/12

OpenStudy (anonymous):

can u please show your working

OpenStudy (anonymous):

well you need to find the greatest common factor of the two denomonators

OpenStudy (anonymous):

which is 12

OpenStudy (gorv):

let teacher =x women = x* 3/4 =3x/4

OpenStudy (gorv):

from 3/4 women 5/6 are married

OpenStudy (anonymous):

so that would be \[9/12 and 10/12\]

OpenStudy (gorv):

\[\frac{ 3 }{ 4 } * \frac{ 5 }{ 6 }\]

OpenStudy (anonymous):

wait i am completely off never mind gorv take it away

OpenStudy (anonymous):

@domisreallytall @gorv thank you :D

OpenStudy (anonymous):

np